OpenStudy (anonymous):

Let f(x)=1/2x, g(x)= 3x^{2} +7, and h(x)= 2x^{3}. Then (f ○ g ○ h) (6) =.

OpenStudy (anonymous):

Ahm. I have small problem solving it because i'm not sure i'm reading it right \[ f(x) = \frac{1}{2x} \\ g(x) = 3x^2 + 7 \\ h(x) = 2x^3 \] all good?

OpenStudy (anonymous):

yes, that is correct.

OpenStudy (anonymous):

\[ (f \circ g \circ h)(x) = f\bigg[\;g\big[\;h(x)\;\big] \bigg] \\ h(6) = 2 \cdot6^3 = 2 \cdot 216 = 432 \\ (g \circ h)(6) = 3 \cdot 432^2 + 7 = 3 \cdot 186624 + 7 = 559879 \\ (f \circ g \circ h)(6) = \frac{1}{2 \cdot 559879} = 0.000000893 \] If i didn't miscalculate :|

OpenStudy (anonymous):

Than you so much!!!